### v1

# Theory-agnostic searches for non-gravitational modes in black hole ringdown

Abstract
In any extension of General Relativity (GR), extra fundamental degrees of freedom couple to gravity. Besides deforming GR forecasts in a theory-dependent way, this coupling generically introduces extra modes in the gravitational-wave signal. We propose a novel theory-agnostic test of gravity to search for these nongravitational modes in black hole merger ringdown signals. To leading order in the GR deviations, their frequencies and damping times match those of a test scalar or vector field in a Kerr background, with only amplitudes and phases as free parameters. This test will be highly valuable for future detectors, which will achieve signal-to-noise ratios higher than 100 (and as high as 1000 for space-based detectors such as LISA). Such sensitivity will allow measurement of these modes with amplitude ratios as low as 0.05 for ground-based detectors (and as low as 0.008 for LISA), relative to the fundamental mode, enabling stringent agnostic constraints or detection of scalar/vector modes. By applying this test to GW150914, GW190521, and GW200129, we find that the current evidence for an extra mode is comparable to that for the first gravitational overtone, but its inclusion modifies the inferred remnant spin.
Introduction. The black hole (BH) spectroscopy program [1, 2, 3, 4] plays a prominent role in the landscape of strong-field tests of General Relativity (GR) [5, 6, 7] and provides a unique method for examining the nature of compact remnants formed post-coalescence [8]. This program focuses on extracting the remnant quasinormal modes (QNMs) [9, 10, 11, 12] during the ringdown phase of a binary merger. In the context of linear perturbation theory, the signal $h(t)$ at intermediate times after the merger is represented by a superposition of the QNMs of the remnant [13], as it transitions towards a stationary configuration. Schematically,
$h(t)=\sum_{i}A_{i}\cos\left(2\pi f_{i}t+\phi_{i}\right)e^{-\frac{t}{\tau_{i}}}\,,$ (1)
where $A_{i}$, $\phi_{i}$, $f_{i}$, $\tau_{i}$ are the amplitude, phase, frequency, and damping time of the $i$-th QNM. If the remnant is a BH, GR predicts that the infinite spectrum of QNMs is uniquely determined by its mass and spin $(M_{f},\chi_{f})$. This provides opportunities for conducting multiple null-hypothesis tests of gravity [14, 15] and investigating the nature of the remnant [16, 17, 18].
As suggested by Lovelock’s theorem [19, 20], an almost unavoidable ingredient of theories beyond GR is the presence of extra degrees of freedom nonminimally coupled to gravity [21, 6]. Examples are ubiquitous and include scalar fields in scalar-tensor theories and Horndeski’s gravity [22] (and their vector counterpart [23]), high-curvature corrections to GR that predict extra (pseudo)scalars and dilaton fields [24, 25, 26], Einstein-Aether [27] and Hořava–Lifshitz [28] gravity that postulate an extra timelike vector field, and massive gravity [29, 30] with both scalar and vector dynamical degrees of freedom (see [6] for a review of GR extensions and their field content). Effective extra degrees of freedom are also unavoidable in any approach that treats GR as the leading order term in an effective-field-theory expansion (e.g., [26]) and in low-energy effective string theories. These nonminimally coupled fields may modify the stationary BH solutions, leading to deviations from the Kerr metric, and/or modify the dynamics of the theory. In either case, two generic predictions are: i) a deformation of the Kerr QNMs,
$f_{i}=f_{i}^{\rm Kerr}(1+\delta f_{i})\,,\qquad\tau_{i}=\tau_{i}^{\rm Kerr}(1+\delta\tau_{i})\,,$ (2)
and ii) the existence of extra modes in the gravitational signal, that can be excited during the ringdown. This second option is due to the fact that the nonminimal coupling between new degrees of freedom and gravity results in coupled systems of linear perturbation equations, which act as a coupled set of oscillators [31, 32, 33]. It is therefore natural to split the ringdown signal (1) into two contributions,
$h(t)=$ $\sum_{i}A_{i}\cos\left(2\pi f_{i}^{\rm Kerr}(1+\delta f_{i})t+\phi_{i}\right)e^{-\frac{t}{\tau_{i}^{\rm Kerr}(1+\delta\tau_{i})}}$
$+\sum_{i}\hat{A}_{i}\cos\left(2\pi\hat{f}_{i}t+\hat{\phi}_{i}\right)e^{-{t}/{\hat{\tau}_{i}}}\,,$ (3)
where we use the hat to denote quantities related to the extra modes.
Standard tests of gravity based on the ringdown are rooted in the first line of Eq. (Theory-agnostic searches for non-gravitational modes in black hole ringdown). Namely, they are aimed at measuring $\delta f_{i}$ and $\delta\tau_{i}$ and check whether they are compatible with the null hypothesis. This can be accomplished in two complementary ways, either through a theory-agnostic or theory-dependent method. However, both approaches have their own limitations. In the theory-agnostic approach one uses the parametrization (2), where $f_{i}^{\rm Kerr}$ and $\tau_{i}^{\rm Kerr}$ are known functions of the BH mass and spin, while $\delta f_{i}$ and $\delta\tau_{i}$ depend on the mass, spin, and all extra fundamental coupling constants of the theory. It is convenient (and natural, from an effective-field-theory perspective) to restrict to small GR deviations that continuously deform the Kerr result, in which case $\delta f_{i},\delta\tau_{i}\ll 1$ and are proportional to some combination of the mass and coupling constants. However, even in this case they are still generic functions of the spin $\chi_{f}$. Current parametrizations either neglect such spin dependence [5, 34, 35] or consider a small spin expansion of each deviation [36, 37, 38], which inevitably inflates the number of free parameters in the model. In the theory-dependent approach, the QNMs are computed in a given theory of gravity. For most theories this can be done again only in the small-coupling limit and very often perturbatively in the spin [39, 40, 31, 41, 42, 43, 44, 45, 46, 47, 48]. To reach good convergence, one needs to push the spin expansion to very high order [47], which is very challenging from the technical point of view. Alternatively, one needs to solve intricate systems of coupled partial differential equations [49, 50, 51, 52, 53, 54]. This approach has the benefit of limiting the number of free parameters to the sole coupling constants and BH spin, but must be performed on a case-by-case basis for every given theory.
Given the above limitations, it would be highly desirable to develop complementary ringdown tests which are both theory-agnostic and accurate. In this work we explore a currently unbeaten path, related to the second line of Eq. (Theory-agnostic searches for non-gravitational modes in black hole ringdown). Namely, we propose to look for extra modes in the ringdown signal, which are not related to deformations of the Kerr ones. For concreteness, let us consider the case of an extra scalar degree of freedom nonminimally coupled to gravity (the same argument applies to other types of fields). Due to the coupling, the gravitational perturbations will contain also scalar modes (e.g., [40, 31, 55] for two concrete examples in theories with quadratic curvature terms),
$\hat{f}_{i}=f_{i}^{{\rm Kerr},\,s=0}(1+\delta\hat{f}_{i})\,,\qquad\hat{\tau}_{i}=\tau_{i}^{{\rm Kerr},\,s=0}(1+\delta\hat{\tau}_{i})\,,$ (4)
where $f_{i}^{{\rm Kerr},\,s=0}$ and $\tau_{i}^{{\rm Kerr},\,s=0}$ are the QNMs of a test scalar field in the Kerr metric, and also in this case $\delta\hat{f}_{i}$ and $\delta\hat{\tau}_{i}$ are complicated, theory-dependent, functions of the mass, spin, and coupling constants. Crucially, in this case the amplitudes $\hat{A}_{i}$ of these modes are proportional to (powers of) the coupling constants [40, 31, 55], and they must vanish in the GR limit. Therefore, to leading order in the corrections, we can neglect $\delta\hat{f}_{i}$ and $\delta\hat{\tau}_{i}$, so that the GR deviations are generically parametrized only by the amplitude of the test-field modes. This is precisely what happens in so-called dynamical Chern-Simons gravity [25, 31], although it is a generic feature.
The above considerations suggest a novel ringdown test of gravity based on the following waveform model
$h(t)=$ $\sum_{i}A_{i}\cos\left(2\pi f_{i}^{\rm Kerr}t+\phi_{i}\right)e^{-\frac{t}{\tau_{i}^{\rm Kerr}}}$
$+\sum_{i}\hat{A}_{i}\cos\left(2\pi f_{i}^{{\rm Kerr},\,s=0}t+\hat{\phi}_{i}\right)e^{-\frac{t}{\tau_{i}^{{\rm Kerr},\,s=0}}}$ (5)
where for simplicity we have neglected the GR deviations in the first line, since those are very well studied by standard ringdown tests (and subjected to the aforementioned limitations). In practice, here we will focus on a standard GR ringdown waveform (first line of Eq. (Theory-agnostic searches for non-gravitational modes in black hole ringdown)) augmented by new extra modes. Remarkably, to leading order these extra modes are known functions of the BH mass and spin, since they are those of a free test (scalar, vector, etc) field propagating on the Kerr metric (see, e.g., [56, 57] for tabulated values). This allows searching for extra modes in a theory-agnostic way, where the amplitudes and phases of the extra modes are the only beyond-GR parameters. In this sense, this test is reminiscent of searches for extra (scalar, vector) polarizations in GW signals in a theory-agnostic fashion [58, 59, 60, 5] and is complementary to ordinary ringdown tests (see, e.g., [61, 5, 34, 35, 62, 63]).
Searching for extra ringdown modes. The ringdown signal comprises of two polarizations and the modes are decomposed in a basis of spin-weighted spheroidal harmonics that depend on the remnant spin inclination angle [63]. Each mode is identified by three integers, $i\equiv(lmn)$, namely the multipolar, azimuthal, and overtone index, respectively, where $n=0$ corresponds to the fundamental tone and $(lmn)=(220)$ is the dominant mode in a non-precessing quasicircular coalescence. For concreteness, here we focus on the most interesting quadrupolar ($l=2$) case, but our test can be applied also to higher-order modes. In particular, our ringdown waveform model has the following parameters:
$\underline{\theta}=\{M_{f},\chi_{f},A_{22j},\phi_{22j},\hat{A}_{220}^{s=0,1},\hat{\phi}_{220}^{s=0,1}\}$ (6)
where $j={0,\dots,N}$, with $N$ the total number of overtones, whereas $\hat{A}_{220}^{s=0,1}$ and $\phi_{220}^{s=0,1}$ are the amplitude and phase of the extra scalar ($s=0$) or vector ($s=1$) $(220)$ mode. We will dub a model with $N$ tones and no extra mode GRN, whereas we will denote as GRN+S (GRN+V) a model with $N$ gravitational tones and an extra scalar (vector) mode. In practice, we define the amplitude ratio $A_{R}$ of a given mode relative to the gravitational fundamental one. We fix the luminosity distance $d_{L}$, the sky location of the system, and the inclination angle. Note that, for small-redshift sources (including those detected so far and presumably the loudest events detected in the future) $d_{L}$ is degenerate with the overall ringdown amplitude and can therefore be neglected without loss of generality. However, the possibility of neglecting the sky location and inclination angle is a prerogative of our test, since it involves only $(l=m=2)$ modes, which have the same pattern functions and spheroidal-harmonic decomposition. Thus, these extrinsic parameters factor out in the waveform, which is an important simplification compared to ordinary tests based on subleading modes with different values of $(l,m)$.
The frequencies and damping times of the relevant modes are shown in the appendix. At variance with overtones, the frequency of the $(220)$ scalar or vector mode is always well separated from that of the fundamental gravitational mode (and hence more easily resolvable from the latter), while the damping time is comparable (and hence the mode survives longer than overtones in the signal, almost as long as the fundamental gravitational mode).
Bayesian analysis on real data. We exemplify our test on real events by performing a Bayesian parameter estimation using the PyCBC Inference code infrastructure [64]. The analysis aims to compute the posterior distribution of the parameters (6). We apply this test to three events: (i) GW150914 [65], the first GW event ever detected by LIGO (and still so far the one with the largest ringdown signal-to-noise ratio (SNR)), for which some debated evidence of overtones has been reported [14, 66, 67, 68, 69, 70, 71, 72, 73, 34, 63]; (ii) GW190521 [74], a peculiar event in the upper mass gap for which a tentative detection of the $(330)$ has been obtained [75, 76] and which is prone also to ringdown amplitude-phase consistency tests [62]; (iii) GW200129, a peculiar loud event showing some tension with GR in some inspiral-merger-ringdown tests [5], tentatively ascribed to mismodelling of precession [77, 78]. For the luminosity distance, inclination, and sky location we adopt the maximum likelihood values reported in [79].
We use a gated-and-inpainted Gaussian likelihood noise model [80, 81, 75] to remove the influence of the pre-peak/non-ringdown times. The strain data within a time interval $t\in[t_{c}+t_{\rm offset}-0.5s,t_{c}+t_{\rm offset}]$ are replaced/inpainted such that the filtered inverse power spectral density is zero at all the times corresponding to the chosen interval [81]. Here, $t_{c}$ is the coalescence time, while $t_{\rm offset}$ defines the time in which we start our ringdown analysis.
Figure 1: $\log_{10}$ Bayes factors for various ringdown models with extra scalar or vector modes with respect to the GR1 model as a function of the offset time $t_{\rm offset}$. Different colors and line styles denote different events, while different markers show the chosen model.
Figure 2: Left panel: Minimum SNR necessary for detecting a scalar mode at $1-\sigma$ confidence level, according to the detectability criterion outlined in [82] using a Fisher matrix approximation. We take as a reference an event with remnant mass $M_{f}=62M_{\odot}$ for a range of amplitude ratios $A_{R,220}\in[0.02,1]$. The blue and red dashed curves denote a different remnant spin $\chi_{f}$ with the phase difference between the GR and scalar modes fixed to $\delta\phi=0$. The dependence on $\delta\phi\in[0,2\pi]$ is bracketed by the corresponding colored bands. The horizontal shaded bands represent the typical ringdown SNR expected for a GW150914-like event observed by the LVK network at design sensitivity (light blue), A+ (light green), ground-based third generation observatories as ET/CE (yellow), and LISA (orange). Finally, the blue dots denote the results of a Bayesian analysis obtained from injecting the scalar ringdown into zero noise for $\chi_{f}=0$. Right: The same quantity shown in a contour plot on the $(A_{R,220},\delta\phi)$ plane for fixed $\chi_{f}=0.67$.
Our main results are summarized in Fig. 1, presenting the statistical evidence for different waveform models for these events and for different choices of $t_{\rm offset}$. We show $\log_{10}{\cal B}_{\texttt{GR1}}^{\rm Model}$, where the Bayes factor ${\cal B}_{\texttt{GR1}}^{\rm Model}$ is the ratio between the evidence of a given model and that of GR1 (i.e., a model containing only the fundamental gravitational mode and the first overtone). According to Jeffreys’ scale criterion [83], a $\log_{10}$ Bayes factor larger than $1$ (resp., $2$) would imply a strong (resp., decisive) Bayesian evidence in favor of a given model relative to GR1. The small values of $\log_{10}{\cal B}_{\texttt{GR1}}^{\rm Model}$ shown in Fig. 1 for any $t_{\rm offset}$ indicate that all models with extra scalar or vector modes have the same evidence as the GR1 one, presumably because the SNR in the ringdown for these events is not sufficiently high to exclude the presence of an extra scalar mode. This is consistent with what we shall discuss below with synthetic data.
Model $M_{f}(M_{\odot})$ $\chi_{f}$ ${A_{220}}\times{10^{20}}$ $A_{R,220}$
GR1 $58^{+27}_{-22}$ $0.35^{+0.49}_{-1.15}$ $0.72^{+0.98}_{-0.50}$
GR0+S $52^{+35}_{-16}$ $-0.18^{+0.94}_{-0.69}$ $0.69^{+1.32}_{-0.57}$ $0.80^{+2.74}_{-0.64}$
GR1+S $52^{+33}_{-16}$ $-0.11^{+0.84}_{-0.76}$ $0.72^{+1.30}_{-0.60}$ $1.95^{+1.85}_{-1.75}$
Table 1: 90% credible intervals for some of the parameters of event GW150914, assuming $t_{\rm offset}=2\,{\rm ms}$ (see Appendix for the posterior distributions). We denote the amplitude ratio of the scalar-to-tensor mode as $A_{R,220}=\hat{A}_{220}^{s=0}/A_{220}$.
Interestingly, despite Fig. 1 showing that there is no statistical evidence for an extra scalar or vector mode, its inclusion affects the posterior distributions of the parameters. This is shown in Table 1 for a representative example of GW150914 analyzed with an extra scalar mode. Other events, different time offsets, or the vector case show qualitatively similar results. While the presence of an extra mode does not affect the distribution of the remnant mass significantly, it contributes to broaden up that of the spin towards smaller values (see Appendix for the posterior distributions). This generic feature can be understood from the fact that an extra $(220)$ scalar/vector mode has a damping time comparable to the fundamental gravitational mode (and the damping time is only mildly sensitive to the spin for $\chi_{f}\lesssim 0.8$) and also has a higher frequency than the $(220)$ and $(221)$ gravitational modes. Thus, interpreting the overtone frequency with an extra scalar/vector mode requires a smaller remnant spin. Also, the amplitude of the fundamental gravitational mode is affected by the extra mode: in the GR0+S and GR1+S models the peak of the $A_{220}$ distribution is smaller because part of the information is contained in the scalar mode. Consequently, the amplitude ratio $A_{R,220}$ between the $(220)$ scalar mode and the ($220$) gravitational mode peaks at some nonzero value.
Forecasts with future observations. In Fig. 2, we forecast the constraining power of this test by computing the minimum ringdown SNR, $\rho_{\rm det}$, for detectability of an extra scalar mode in the GR0+S model, for different amplitude ratios, remnant spins, and phase differences $\delta\phi=\phi_{220}-\hat{\phi}_{220}^{s=0}$. (Hereafter we focus on the scalar case, since the vector case gives qualitatively similar results, see Appendix.) $\rho_{\rm det}$ is defined following Refs. [82, 84, 85], namely as the SNR such that $\sigma_{A_{R}}=A_{R,220}$, where the statistical error $\sigma_{A_{R}}$ has been computed with a fully numerical Fisher information matrix [84, 85]. Note that, in contrast to overtones [68, 82], resolving the scalar mode is relatively easy since the frequency is significantly different from the gravitational one. For this reason, $\rho_{\rm det}$ is always larger than the SNR threshold required to resolve the extra scalar mode.
The left panel of Fig. 2 shows that $\rho_{\rm det}$ decreases monotonically as $\chi_{f}$ increases – so highly spinning remnants favor this test – and that the phase difference $\delta\phi$ has a negligible impact for small amplitude ratios, while it can significantly affect the SNR at large amplitudes, especially for slowly-spinning remnants. The dependence on $\delta\phi$ can be better appreciated from the right panel of Fig. 2, showing a contour plot of the two-dimensional function $\rho_{\rm det}(A_{R,220},\delta\phi)$ for $\chi_{f}=0.67$. In the region of small amplitude ratios, the minimum SNR has a very simple scaling
$\rho_{\rm det}\approx 9.3\frac{g(\chi_{f},\delta\phi,A_{R,220})}{A_{R,220}}\,.$ (7)
where $g\approx 1+0.56A_{R,220}-0.32\chi_{f}+0.14\cos(\delta\phi+0.98)$, and for $A_{R,220}\leq 0.12$ and $\chi_{f}\leq 0.9$ the fit is accurate within $10\%$.
In the left panel of Fig. 2 we also provide reference values of the ringdown SNR for current and future detectors. In particular, third-generation ground-based detectors [86] such as Cosmic Explorer [87, 88, 89] and the Einstein Telescope [90, 91, 92] are expected to observe a few events per year with ringdown SNR greater than 100 [85]. Likewise, GW space interferometers such as LISA [93] are expected to detect up to a dozen of massive BH mergers with ringdown SNR greater than 1000, depending on the massive BH population [94, 84]. Our results show that a ringdown SNR of 150 (resp., 1000) would yield a constraint $A_{R,220}\lesssim 0.05$ (resp., 0.008) for $\chi_{f}\approx 0.9$, with only mild dependence on $\delta\phi$. Very similar results apply to the detectability of an extra vector mode.
This plot also confirms that the constraining power of the test is very limited when the ringdown SNR is around 10, which is a rough and even optimistic estimate for the previously analyzed events GW150914, GW190521, and GW200129.
While these results were obtained using a Fisher-matrix approximation to explore the entire parameter space, in the left panel of Fig. 2 we also compare them with results obtained from synthetic injections at zero noise using the same Bayesian analysis discussed above for real-data events. The results of the full parameter estimation are in good agreement with the Fisher approximation, as expected given the large SNR.
Discussion. Any theory predicting extra ringdown modes should presumably predict also deviations from the standard gravitational Kerr QNMs, in which case one could argue that GW detectors are more sensitive to phase differences (and hence to QNM shifts) rather than amplitude differences, so that our test could have less constraining power than ordinary ringdown tests. However, there are known examples of theories predicting zero or negligible QNM shifts but extra modes, in which case our test can be superior to ordinary BH spectroscopy. An example is dynamical Chern-Simons gravity, wherein for a Schwarzschild BH the GR QNMs are unchanged, but the ordinary ringdown contains also extra scalar modes [31]. In the Kerr case also the GR QNMs are modified, but the deviations are suppressed by powers of the BH spin [44], so the convenience of our method will likely depend on the remnant’s spin.
As future extensions, it would be interesting to consider a specific theory and compare the constraints on the coupling constant(s) placed by our test with those of ordinary BH spectroscopy. This would require estimates of both ordinary QNM shifts and excitation amplitudes of extra modes in a given theory, both of which have recently become available for some theories [46, 47, 48, 95, 96, 97, 98, 99, 100, 101, 102]. Another possible avenue of exploration is to use the recently introduced QNM filtering technique [73, 35, 34] to search for extra (scalar, vector, etc) modes.
Acknowledgements. We would like to thank Collin Capano and Alexander Nitz for clarifications on PyCBC inference, and Gabriele Franciolini for insightful comments. This work is partially supported by the MUR PRIN Grant 2020KR4KN2 “String Theory as a bridge between Gauge Theories and Quantum Gravity”, by the FARE programme (GW-NEXT, CUP: B84I20000100001), and by the INFN TEONGRAV initiative. Some numerical computations have been performed at the Vera cluster supported by the Italian Ministry of Research and by Sapienza University of Rome, and with the University of Birmingham’s BlueBEAR HPC service.
Appendix A Supplemental results
Here we present some results complementing those discussed in the main text.
In Fig. 3 we show the frequencies and damping times of some representative QNMs, including scalar, vector, and gravitational modes as functions of the remnant’s spin. Note that the frequency of the $(220)$ scalar or vector mode is always well separated from that of the fundamental gravitational mode, while its damping time is comparable. These are advantages with respect to overtones, which are instead harder to resolve and decay more rapidly.
Figure 3: Frequencies (left) and damping times (right) for a few gravitational modes and for $(220)$ scalar and vector modes of Kerr as a function of the spin. Results are normalized for a remnant mass compatible with GW150914, $M_{f}=62\,M_{\odot}$.
Figure 4 presents an example of posterior distributions of some waveform parameters obtained from our Bayesian analysis on real data, namely the case of GW150914 analyzed with various models. Although, as discussed in the main text, there is no statistical evidence for extra scalar modes in the data, their inclusion in the parameter estimation affects the posterior of $\chi_{f}$ and $A_{220}$.
[…]
Finally, in Fig. 5, we compare some representative posterior distributions obtained from the Bayesian inference on real data with forecasts using injections at higher SNR (equal to 100). We consider a ringdown model with an extra scalar mode (GR0+S, left panel) and with an extra vector mode (GR0+V, right panel). For the injection simulations, we inject the values corresponding to the maximum likelihood values of the real data. We notice that, at higher SNRs, the precision of the parameters’ distribution improves, and the signal is well reconstructed. This can be easily observed with the distributions of $\chi_{f}$, which in the case of the real data is spread, due to the effect of the gated Gaussian noise. Also in this case we note that the results for the scalar or vector case are very similar.
Figure 5: Posterior distributions for real data (magenta) vs injection+recovery (blue) parameter estimation simulations with model GR0+S (left) and GR0+V (right). The levels of the 2D distributions indicate the 68% and the 90% credible intervals, while the black lines correspond to the injected values. Note that, in this case, we indicate the fundamental amplitude $A_{220}$ rescaled by $d_{L}/M^{(m)}_{f}$, where $M^{(m)}_{f}$ is the final mass of the remnant expressed in meters.